The following data were obtained in an enzyme study: [S] (HM) v (HM mini) 2.91 2.00 3.00 3.83 4.00 4.35 4.87 5.00 6.00 5.42 7.00 5.75 6.21 8.00 9.00 6.36 10.00 6.73 Use these data to construct a Michaelis-Menten plot, a Lineweaver-Burk plot, and an Eadie-Hofstee plot. Fit the data in the L-B and E-H plots and use the fit parameters to determine KM and vs in each case. Comment on how well the KMs and vss agree with each other.
Answer #1

[S] µM

v, µMmin-1

1/[S] µM-1

1/v, min/µM

v/[s] min-1

0

0

-

-

-

2

2.91

0.50

0.34

1.46

3

3.83

0.33

0.26

1.28

4

4.35

0.25

0.23

1.09

5

4.87

0.20

0.21

0.97

6

5.42

0.17

0.18

0.90

7

5.75

0.14

0.17

0.82

8

6.21

0.13

0.16

0.78

9

6.36

0.11

0.16

0.71

10

6.73

0.10

0.15

0.67

a) Michaelis-Menten Plot: Plot of Reaction rate, v vs substrate concentration, [S]

Michaelis-Menten Plot 10 [S], HM

b) Lineweaver-Burk Plot: Plot of 1/v vs 1/[S]

Lineweaver-Burk Plot 0.40 0.35 0.30 0.25 2 0.20 y 0.4817x+0.104 0.15 0.10 0.05 0.00 0.00 0.10 0.20 0.30 0.40 0.50 0.60

y-intercept = 0.104 = 1/Vs

Vs = 9.6 µMmin-1

Slope = 0.4817 = Km/Vs

Km = 0.4817 * 9.6 = 4.63 µM

c) Eadie-Hoftee Plot: Plot of v vs v/[S]

Eadie-Hofstee Plot y:-4.7599X+9.7462m 0.60 0.80 1.00 1.20 1.40 1.60 v/[S], 1/min

y-intercept = 9.75 = Vs

Vs = 9.75 µMmin-1

Slope = -4.76 = - Km

Km = 4.76 µM

*Results of Lineweaver-Burk and Eadie-Hoftee are close in value (<0.2 – 0.3 % difference).